Constants Part A A 235-g sample of a substance is heated to 320 °C and then plunged into a 105-g aluminum calorimeter cup containing 195 g of water and a 17-g glass thermometer at 10.5 ° C. The final temperature is 35.0° C. The value of specific heat for aluminium is 900 J/kg-C ,for glass is 840 J/kg . Co , and for water is 4186 J/kg-C . What is the specific heat of the substance? (Assume no water boils away.) Express your answer using three significant figures ? 938.390 SubmitPrev X Incorrect, Try Again; 5 attempts remaining ious Answers Request Answer Provide Feedback Next >
Answer #1

here,

mass of substance , m1 = 235 g = 0.235 kg

mass of alumunium , m2 = 105 g = 0.105 kg

mass of water , m3 = 195 g = 0.195 kg

mass of glass , m4 = 17 g = 0.017 kg

let the specifc heat of substance be Cs

using conservation of energy

heat lost by substance = heat gained by (alumunium + water + glass)

m1 * Cs * ( 320 - 35) = m2 * Ca * ( 35 - 10.5) + m3 * Cw * ( 35 - 10.5) + m4 * Cg * ( 35 - 10.5)

0.235 * Cs * ( 320 - 35) = 0.105 * 900 * ( 35 - 10.5) + 0.195 * 4186 * ( 35 - 10.5) + 0.017 * 840 * ( 35 - 10.5)

solving for Cs

Cs = 338.4 J/kg C

the specifc heat capacity of substance is 338.4 J/kg C